eight roses grew in a garden 1/4 of them were picked and put in a vase how many were left

Answer by stanbon(75887)

eight roses grew in a garden 1/4 of them were picked and put in a vase how many were left
-----
# left = 8 - (1/4)8
----
# left = 8-2 = 6
=============================
Cheers,
Stan H.